Suppose that grade point averages of undergraduate students at one university have a bell-shaped distribution with a mean of 2.51 and a standard deviation of 0.43. Using the empirical rule, what percentage of the students have grade point averages that are no more than 3.37? Please do not round your answer.
Answer by ewatrrr(24785)

3.37 = 2.51 + 2*.43
3.37 is 2 SD Above or to the right of the mean
percentage of the students have grade point averages that are no more than 3.37:
.50 + 47.7 = 97.7
one standard deviation from the mean accounts for about 68.2% of the set
two standard deviations from the mean account for about 95.4% *******95.4/2 = 47.7
and three standard deviations from the mean account for about 99.7%.
Important to Understand z -values as they relate to the Standard Normal curve: